If 3 balls are drawn at random from a bag containing 3 red and 4 blue balls, what is the expected number of red balls in the sample?
Answer by stanbon(48568)

If 3 balls are drawn at random from a bag containing 3 red and 4 blue balls, what is the expected number of red balls in the sample?
----------
Random variable values are 0,1,2,3
Associated probabilities are:
P(0 red)= 4C3/7C3=4/[7*6*5]/[1*2*3]=0.11
P(1 red)= [3C1*4C2/7C3]=0.51
P(2 red)= [3C2*4C1]/7C3]=0.34
P(3 red)= [3C3]/7C3=0.03
----------------
Expected number of red = 0.11*0+0.51*1+0.34*2+0.03*3=1.3 red apples
----------
Cheers,
Stan H.